A sample of nitrogen gas was collected via water displacement. Since the nitrogen was collected via water displacement, the sample is saturated with water vapor. If the total pressure of the mixture at 21°C is 1.73 atm, what is the partial pressure of nitrogen? The vapor pressure of water at 21°C is 18.7mm Hg.
1
Expert's answer
2019-11-20T06:54:20-0500

P(sum)=P(water)+P(N2);

18.7 mm Hg=0.02 atm

P(N2)=P(sum)-P(water)=1.73-0.02=1.71 atm